OpenStudy (anonymous):

Solve each system by graphing y = -4x y = - x + 6 (0, 6) (-6, 6) (-2, 8) (-8, 0)

OpenStudy (anonymous):

(-3, -4)

OpenStudy (anonymous):

Blue line is (x-y=1) ans green line is (x+y=7)

OpenStudy (anonymous):

(1, 3) is the solution... blue line is 6x + 2y = 12 and green line is y = 3x